OpenStudy (anonymous):

PLEASE HELP solve the inequality. show your work |r+3|≥ 7 i have the answer r≥ 4

OpenStudy (texaschic101):

| r + 3 | >= 7 (r + 3) >= 7 r >= 7 - 3 r >= 4 (r + 3) <= -7 r <= -7 - 3 r <= -10 so r >= 4 and r <= -10 this will help.. https://www.purplemath.com/modules/absineq.htm

OpenStudy (anonymous):

so theres two answers?